# Primordial Black Hole Reformation in the Early Universe

### v1

@@ V Gravitational Wave Signal
The rapid evaporation and reheating after the eMD phase leads to a coincident GW signal. The stochastic GW background emitted during the PBH reheating was computed in Refs. Inomata et al. (2020); Papanikolaou et al. (2021); Domènech et al. (2021a, b), and showed that the rapid reheating characteristic of Hawking evaporation led to significant tensor mode production. The resulting GW spectrum is highly wavenumber dependent, $\Omega_{\rm GW}\sim k^{11/3}$ and peaks at the ultraviolet cutoff $k_{\rm UV}$ Domènech et al. (2021a),
$\begin{split}\Omega_{\rm GW,0}(k_{\rm UV})\simeq&2.1\times 10^{-10}\left(\frac{k_{\rm UV}}{k_{\rm RH}}\right)^{17/3}\left(\frac{k_{\rm eq}}{k_{\rm UV}}\right)^{8}\\ \simeq&3.2\times 10^{11}\beta_{\rm if}^{16/3}\left(\frac{\gamma}{0.5}\right)^{8/3}\left(\frac{M_{\rm PBH}}{1\,\text{g}}\right)^{34/9}\leavevmode\nobreak\ .\end{split}$ (18)
Depending on the parameters, the stochastic GW background generated by PBH reheating may be detected by planned upgrades to aLIGO Abbott et al. (2020) as well as upcoming experiments such as LISA Amaro-Seoane et al. (2017), Einstein Telescope Hild et al. (2011), and Cosmic Explorer Srivastava et al. (2022). Additionally, the emission of GWs can be constrained by the current bound on $\Delta N_{\rm eff}\lesssim 0.5$ by Planck TT+low E observations Aghanim et al. (2020) and similarly by BBN observations Arbey et al. (2021), and be potentially detected by the CMB-S4 experiment which has a sensitivity limit of $\Delta N_{\rm eff}=0.06$ Abazajian et al. (2019).
Figure 3: The coincident GW signature of two models of reformed PBH populations. The initial populations of extremely light PBHs result in a high frequency stochastic GW background after evaporation, higher than aLIGO or ET’s reachable range Schmitz (2021). These GWs contribute to $\Delta N_{\rm eff}$ at BBN Caprini and Figueroa (2018) and could be detected by the upcoming CMB-S4 experiment Abazajian et al. (2019).
@@ VI PBH Reformation after Reheating
After reheating, the inhomogenous PBH distribution is converted into radiation inhomogeneities. Horizon-sized overdense regions could then grow and recollapse into a PBH Carr (1975). We estimate the abundance of these reformed PBHs during radiation domination using the formula Young et al. (2014); Carr and Kuhnel (2020)
$\beta_{\rm rf}=\gamma\leavevmode\nobreak\ \textrm{erfc}\left(\frac{\delta_{c}}{\sqrt{2}\sigma_{\delta}}\right)\leavevmode\nobreak\ ,$ (19)
where the fiducial value for the critical collapse threshold is taken to be $\delta_{c}=0.45$ Musco and Miller (2013); Carr and Kuhnel (2020) and the variance is evaluated at reheating $\sigma_{\delta}(k_{\rm RH}^{-1},t_{\rm RH})$. Due to the erfc in Eq. (19), the abundance of reformed PBHs depends sensitively on the ratio $\delta_{c}/\sigma_{\delta}$, with significant production occurring only for $\sigma\gtrsim\mathcal{O}(10^{-2})$.
This makes any significant reformation after reheating practically rely on the clustering of the initial PBHs. Here we follow the parameterization of Ref. Domènech et al. (2021a), which is convenient for both the computation of the GW signal and PBH reformation,
$P_{\delta,\rm cl}(k)=P_{\delta}(k)\left(\frac{k}{k_{\rm UV}}\right)^{3}\left(1+F_{\rm cl}\left(\frac{k}{k_{\rm cl}}\right)^{n_{\rm cl}}\right)\leavevmode\nobreak\ ,$ (20)
@@
$\Omega_{\rm GW,cl,0}(k_{\rm UV})=F_{\rm cl}^{2}\left(\frac{k}{k_{\rm cl}}\right)^{2n_{\rm cl}}\Omega_{\rm GW,0}(k_{\rm UV})\leavevmode\nobreak\ .$ (21)
If these clustered PBHs induce the last eMD, right before the standard cosmological history, the current $N_{\rm eff}$ bound results in a constraint on the $M_{\rm if},\beta$, and clustering parameters Domènech et al. (2021a, b),
@@
×$\times\left(\frac{k_{\rm cl}}{k_{\rm UV}}\right)^{3n_{\rm cl}/8}\left(\frac{\gamma}{0.5}\right)^{-1/2}\left(\frac{M_{\rm if}}{1\leavevmode\nobreak\ g}\right)^{-17/24}\leavevmode\nobreak\ .\quad$
Figure 4: Contour plot showing the resulting PBH abundance $\beta_{\rm rf}$ varying the initial PBH density $\beta_{\rm if}$ and the clustering strength $F_{\rm cl}$, with other relevant parameters shown. Significant reformation of PBHs can occur after reheating for $\beta_{\rm if}\lesssim 10^{-2}$ depending on the initial clustering parameters. A second round of eMD could occur, resulting in cascading PBH formation, like the yellow cross representing Case 1 in Table 2.
In Fig. 4, we show the abundance of reformed PBHs, $\beta_{\rm rf}$, as a function of initial abundance $\beta_{\rm if}$ and clustering parameters. For initial PBH densities down to $\beta_{\rm if}\sim 10^{-2}$, significant quantities of reformed PBHs can be produced. These may be sufficient to induce a second eMD era, or even result in cascading production of increasingly heavy reformed PBHs. To avoid interfering with BBN processes, only reformed PBHs with mass $M\lesssim 4\times 10^{8}g$ can undergo a second eMD.
@@
Table 2: Example cases where the PBH reformation after reheating induces 2nd eMD. Both cases result in the reformed PBH population of mass $M_{\rm PBH,rf}\simeq 10^{8}\,\text{g}$ and $\beta_{\rm rf}\simeq 10^{-9}$, being at the boundary of the allowed last PBH eMD (green shaded of Fig. 2).
Table 2 lists the parameters for two example cases that result in reformed PBH populations after reheating of $M_{\rm PBH,rf}\simeq 10^{8}\,\text{g}$ and $\beta_{\rm rf}\simeq 10^{-9}$. Case 1 assumes a typical horizon collapse during the radiation-dominated era right after the inflationary reheating, and Case 2 assumes a fast first-order phase transition at a very high temperature, producing subhorizon scale initial PBHs Hawking et al. (1982); Moss (1994); Jung and Okui (2021). Although the formation temperature $T_{\rm if}$ of Case 1 slightly exceeds the bound on the inflation scale coming from the null detection of tensor modes in CMB observations, these excess GWs could be suppressed by redshift during the eMD phases. For Case 2, the PBHs do not strictly follow the mass-temperature relation, Eq. (1), which is possible in first-order phase transition formation scenarios.
@@ VII Conclusion
A small population of extremely light PBHs with mass $M\lesssim 4\times 10^{8}g$ can engender an eMD era. The growth of structure during matter domination could seed the overdensities that collapse into larger, reformed PBHs. Our computation shows that PBH reformation is applicable to PBHs initially formed in a variety of scenarios, from inflationary fluctuations to first order phase transitions. We explored the possibility of PBH reformation during and right after the eMD era, which could result in evaporating reformed PBHs emitting high energy cosmic rays and gamma ray bursts. These signals could be matched with a coincident GW signature that results from the rapid evaporation and reheating of the initial PBH population. Correlated observations of these two phenomena would be indicative of the PBH reformation mechanism.
@@ Acknowledgements.
The authors thank Donghui Jeong for many helpful discussions. T.H.K. is supported by KIAS Individual Grant PG095201 at Korea Institute for Advanced Study. P.L. is supported by KIAS Individual Grant 6G097701.
@@ Appendix A Numerical Computation of Linear Growth Equations
We list the evolution equations for cosmological perturbations Seljak and Zaldarriaga (1996) in the conformal Newtonian gauge. We have two scalar potentials, the hierarchical moments of radiation perturbation, the PBH density perturbation and velocity. In expressing the evolution, we use $\ln y=\ln(a/a_{\rm eq})$ as a time variable. The two scalar potentials evolve as
$\frac{d\Phi}{d\ln y}=\Psi-\frac{\kappa^{2}}{3}\frac{2y^{2}}{y+1}\Phi+\frac{1}{2}\left(\frac{y}{y+1}\delta_{\rm PBH}+\frac{1}{y+1}4\Theta_{r,0}\right)\leavevmode\nobreak\ ,$ (23)
@@
$\frac{d\Theta_{r,1}}{d\ln y}$ $=$ $\frac{\kappa}{3}\sqrt{\frac{2y^{2}}{y+1}}(\Psi+\Theta_{r,0}-2\Theta_{r,2})\leavevmode\nobreak\ ,$ (26)
$\frac{d\Theta_{r,2}}{d\ln y}$ $=$ $\frac{\kappa}{5}\sqrt{\frac{2y^{2}}{y+1}}(2\Theta_{r,1}-3\Theta_{r,2})\leavevmode\nobreak\ ,$ (27)
@@
with the Boltzmann hierarchy truncated at $l_{\rm max}$ through Ma and Bertschinger (1995)
@@
For superhorizon modes, we have constant values of $\delta_{\rm PBH}=2/5$ and $\Psi=-1/5$ for $y\gg 1$. Since the two have comparable sizes, we use the gauge independent curvature perturbation $\zeta=-\Psi+\delta_{\rm PBH}/3=1/3$. For superhorizon modes, the density perturbation that determines PBH formation is obtained by Young (2019)

### v2

@@ V Gravitational Wave Signal
The rapid evaporation and reheating after the eMD phase leads to a coincident GW signal. The stochastic GW background emitted during the PBH reheating was computed in Refs. [40, 41, 42, 43, 44, 45], and showed that the rapid reheating characteristic of Hawking evaporation led to significant tensor mode production. The resulting GW spectrum is highly wavenumber dependent, $\Omega_{\rm GW}\sim k^{11/3}$ and peaks at the ultraviolet cutoff $k_{\rm UV}$ [42],
$\begin{split}\Omega_{\rm GW,0}(k_{\rm UV})\simeq&\,2.1\times 10^{-10}\left(\frac{k_{\rm UV}}{k_{\rm RH}}\right)^{17/3}\left(\frac{k_{\rm eq}}{k_{\rm UV}}\right)^{8}\\ \simeq&\,3.2\times 10^{11}\beta_{\rm if}^{16/3}\left(\frac{\gamma}{0.5}\right)^{8/3}\left(\frac{M_{\rm PBH}}{1\,\text{g}}\right)^{34/9}~{}.\end{split}$ (17)
Depending on the parameters, the stochastic GW background generated by PBH reheating may be detected by planned upgrades to aLIGO [90] as well as upcoming experiments such as LISA [91], Einstein Telescope [92], and Cosmic Explorer [93]. Additionally, the emission of GWs can be constrained by the current bound on $\Delta N_{\rm eff}\lesssim 0.5$ by Planck TT+low E observations [94] and similarly by BBN observations [95], and be potentially detected by the CMB-S4 experiment which has a sensitivity limit of $\Delta N_{\rm eff}=0.06$ [96].
Figure 3: The coincident GW signature of two models of reformed PBH populations. The initial populations of extremely light PBHs result in a high frequency stochastic GW background after evaporation, higher than aLIGO or ET’s reachable range [97]. These GWs contribute to $\Delta N_{\rm eff}$ at BBN [98] and could be detected by the upcoming CMB-S4 experiment [96].
@@ VI PBH Reformation after Reheating
After reheating, the inhomogenous PBH distribution is converted into radiation inhomogeneities. Horizon-sized overdense regions could then grow and recollapse into a PBH [2]. We estimate the abundance of these reformed PBHs during RD using the formula [51, 56]
$\beta_{\rm rf}=\gamma~{}\textrm{erfc}\left(\frac{\delta_{c}}{\sqrt{2}\sigma_{\delta}}\right)~{},$ (18)
where the fiducial value for the critical collapse threshold is taken to be $\delta_{c}=0.45$ [81, 56] and the variance is evaluated at reheating $\sigma_{\delta}(k_{\rm RH}^{-1},t_{\rm RH})$. Due to the erfc in Eq. (18), the abundance of reformed PBHs depends sensitively on the ratio $\delta_{c}/\sigma_{\delta}$, with significant production occurring only for $\sigma\gtrsim\mathcal{O}(10^{-2})$.
This makes any significant reformation after reheating practically rely on the clustering of the initial PBHs. Here we remain agnostic about the clustering mechanism, and adopt a phenomenological parameterization of Ref. [42], which is convenient for both the computation of the GW signal and PBH reformation,
$P_{\delta,\rm cl}(k)=P_{\delta}(k)\left(\frac{k}{k_{\rm UV}}\right)^{3}\left(1+F_{\rm cl}\left(\frac{k}{k_{\rm cl}}\right)^{n_{\rm cl}}\right)~{},$ (19)
@@
$\Omega_{\rm GW,cl,0}(k_{\rm UV})=F_{\rm cl}^{2}\left(\frac{k}{k_{\rm cl}}\right)^{2n_{\rm cl}}\Omega_{\rm GW,0}(k_{\rm UV})~{}.$ (20)
If these clustered PBHs induce the last eMD, right before the standard cosmological history, the current $N_{\rm eff}$ bound results in a constraint on the $M_{\rm if},\beta$, and clustering parameters [42, 43],
@@
×$\times\left(\frac{k_{\rm cl}}{k_{\rm UV}}\right)^{3n_{\rm cl}/8}\left(\frac{\gamma}{0.5}\right)^{-1/2}\left(\frac{M_{\rm if}}{1~{}\textrm{g}}\right)^{-17/24}~{}.\quad$
Figure 4: Contour plot showing the resulting PBH abundance $\beta_{\rm rf}$ varying the initial PBH density $\beta_{\rm if}$ and the clustering strength $F_{\rm cl}$, with other relevant parameters shown. Significant reformation of PBHs can occur after reheating for $\beta_{\rm if}\gtrsim 10^{-2}$ depending on the initial clustering parameters. A second round of eMD could occur, resulting in cascading PBH formation, like the yellow cross representing Case 1 in Table 2.
In Fig. 4, we show the abundance of reformed PBHs, $\beta_{\rm rf}$, as a function of initial abundance $\beta_{\rm if}$ and clustering parameters. For initial PBH densities down to $\beta_{\rm if}\sim 10^{-2}$, significant quantities of reformed PBHs can be produced. These may be sufficient to induce a second eMD era, or even result in cascading production of increasingly heavy reformed PBHs. To avoid interfering with BBN processes, only reformed PBHs with mass $M\lesssim 4\times 10^{8}\,\text{g}$ can undergo a second eMD.
@@
Table 2: Example cases where the PBH reformation after reheating induces 2nd eMD. Both cases result in the reformed PBH population of mass $M_{\rm PBH,rf}\simeq 10^{8}\,\text{g}$ and $\beta_{\rm rf}\simeq 10^{-9}$, being at the boundary of the allowed last PBH eMD (green shaded region of Fig. 2).
Table 2 lists the parameters for two example cases that result in reformed PBH populations of $M_{\rm PBH,rf}\simeq 10^{8}\,\text{g}$ and $\beta_{\rm rf}\simeq 10^{-9}$. Case 1 assumes a typical horizon collapse during the radiation-dominated era right after the inflationary reheating, and Case 2 assumes a fast first-order phase transition at a very high temperature, producing subhorizon scale initial PBHs [7, 8, 10]. Although the formation temperature $T_{\rm if}$ of Case 1 slightly exceeds the bound on the inflation scale coming from the null detection of tensor modes in CMB observations, these excess GWs could be suppressed by redshift during the eMD phases. For Case 2, the PBHs do not strictly follow the mass-temperature relation, Eq. (1), which is possible in first-order phase transition formation scenarios.
@@ VII Conclusion
A small population of extremely light PBHs with mass $M\lesssim 4\times 10^{8}\textrm{g}$ can engender an eMD era. The growth of structure during matter domination could seed the overdensities that collapse into larger, reformed PBHs. Our computation shows that PBH reformation is applicable to PBHs initially formed in a variety of scenarios, from inflationary fluctuations to first order phase transitions. We explored the possibility of PBH reformation during and right after the eMD era, which could result in evaporating reformed PBHs emitting high energy cosmic rays and gamma ray bursts. These signals could be matched with a coincident GW signature that results from the rapid evaporation and reheating of the initial PBH population. Correlated observations of these two phenomena would be indicative of the PBH reformation mechanism.
@@ Acknowledgements.
The authors thank Donghui Jeong for many helpful discussions and the referee for useful suggestions in improving the paper. T.H.K. is supported by KIAS Individual Grant PG095201 at Korea Institute for Advanced Study. P.L. is supported by KIAS Individual Grant 6G097701.
@@ Appendix A Numerical Computation of Linear Growth Equations
We list the evolution equations for cosmological perturbations [99] in the conformal Newtonian gauge. We have two scalar potentials, the hierarchical moments of radiation perturbation, the PBH density perturbation and velocity. In expressing the evolution, we use $\ln y=\ln(a/a_{\rm eq})$ as a time variable. The two scalar potentials evolve as
$\frac{d\Phi}{d\ln y}=\Psi-\frac{\kappa^{2}}{3}\frac{2y^{2}}{y+1}\Phi+\frac{1}{2}\left(\frac{y}{y+1}\delta_{\rm PBH}+\frac{1}{y+1}4\Theta_{r,0}\right)~{},$ (22)
@@
$\frac{d\Theta_{r,1}}{d\ln y}$ $=$ $\frac{\kappa}{3}\sqrt{\frac{2y^{2}}{y+1}}(\Psi+\Theta_{r,0}-2\Theta_{r,2})~{},$ (25)
$\frac{d\Theta_{r,2}}{d\ln y}$ $=$ $\frac{\kappa}{5}\sqrt{\frac{2y^{2}}{y+1}}(2\Theta_{r,1}-3\Theta_{r,2})~{},$ (26)
@@
with the Boltzmann hierarchy truncated at $l_{\rm max}$ through [100]
@@
For superhorizon modes, we have constant values of $\delta_{\rm PBH}=2/5$ and $\Psi=-1/5$ for $y\gg 1$. Since the two have comparable sizes, we use the gauge independent curvature perturbation $\zeta=-\Psi+\delta_{\rm PBH}/3=1/3$. For superhorizon modes, the density perturbation that determines PBH formation is obtained by [101]